# GAP-LA: GPU-Accelerated Performance-Driven Layer Assignment

### v1

@@ Abstract
Layer assignment is critical for global routing of VLSI circuits. It converts 2D routing paths into 3D routing solutions by determining the proper metal layer for each routing segments to minimize congestion and via count. As different layers have different unit resistance and capacitance, layer assignment also has significant impacts to timing and power. With growing design complexity, it becomes increasingly challenging to simultaneously optimize timing, power, and congestion efficiently. Existing studies are mostly limited to a subset of objectives. In this paper, we propose a GPU-accelerated performance-driven layer assignment framework, GAP-LA, for holistic optimization the aforementioned objectives. Experimental results demonstrate that we can achieve 0.3%-9.9% better worst negative slack (WNS) and 2.0%-5.4% better total negative slack (TNS) while maintaining power and congestion with competitive runtime compared with ISPD 2025 contest winners, especially on designs with up to 12 millions of nets
Index Terms:
Physical design, Global routing, Layer assignment, Timing, Power, GPU acceleration.I Introduction
@@
Delay-driven layer assignment focuses on minimizing maximal or average delay of routing segments under congestion constraints. [12, 13] prioritize nets with large delay and formulate the delay optimization problem as an ILP problem. [14, 15, 16] take coupling effect, slew and non-default-rule routing into account to better handle delay optimization problem in advanced technology node. They also apply edge weighting strategy to optimize maximal and average delay of sinks effectively. It usually takes around ten minutes for those CPU layer assignment frameworks to handle the design with around ten million pins and one million GCells.
@@ II-E Evaluation Metrics
where are input for each design to adjust the weights of different metrics and are input reference metric value. is the number of end points in the design and is total overflow of all the GCell edges. In our evaluation flow, and are reported by open source tool OpenROAD [28]. Note that, value of does not influence the comparison among different solutions because value of is constant. Table I lists the detailed information of 12 cases in ISPD2025 contest.
@@
as its overflow, where is for , for , and denotes the overflow weight of layer , which is given in the benchmark.
@@ III-A Critical Net Analysis-based Net Ordering and Batching Strategy
The delay of critical paths is of paramount importance for design timing closure, and our solution aims to minimize critical path delay as much as possible. However, since layer assignment is a net-based optimization operation, the delay of critical paths cannot be optimized directly. Therefore, identifying the criticality of each net is essential.
Figure 5 depicts an example of a static timing analysis graph with three primary inputs and two primary outputs. The static timing analysis engine reports one critical path for each primary output: Critical path 1 traverses nets , and while critical path 2 traverses nets , and . We define net criticality as the number of critical paths traversing the net. In example shown in Figure 5, net exhibits the highest criticality, whereas nets and have lower criticality, and all other nets have the lowest criticality. Intuitively, optimizing a net with higher criticality yields greater benefits, as it may dominate the arrival times of multiple primary outputs.
@@
However, the strategy described above is purely sequential and must be adapted for parallel computing. Our observations indicate that nets with similar criticality and slack have comparable impacts on final design performance. As a result, strict net ordering can be relaxed: we partition the net set into multiple subsets, enabling parallel assignment within each subset with minimal quality degradation. For critical and semi-critical nets, let denote the maximum criticality. We place nets with criticality into the first subset, into the second subsets, into the third subsets and so on. For other nets with bad slack, we place nets with slack into the first subset, into the second subset, into the third subset and so on. For non-critical nets (better than WNS), we employ a traditional congestion-driven strategy instead, as these nets are far less likely to degrade design performance. Algorithm 1 outlines our critical net analysis-based net ordering and batching strategy flow.
@@ III-B Layer Assignment Directed Tree
To facilitate the extraction of parasitic parameters and the implementation of a fine-grained parallel layer assignment algorithm, we designed a directed tree structure. In our layer assignment algorithm, we assume that a straight 2D routing segment should be assigned to a single layer without layer shifting, which is only performed at bends and Steiner points. To satisfy these constraints, we introduce additional nodes to ensure all inter-node connections are strictly straight, so layer shifting operations are confined to nodes in the directed tree. As illustrated in Figure 6a and Figure 6b, node 6 is added to bridge node 4 and node 7, ensuring no routing paths with bend points exist in the directed tree.
Based on the tree discussed above, we extract parasitic parameters for static timing analysis. The directed tree acts as a subgraph in the STA graph, with nodes connected to relevant pin nodes in the STA graph via zero-resistance edges. For horizontal (vertical) wire connections, we calculate resistance (capacitance) using the average per-unit-length resistance (capacitance) of all horizontal (vertical) wires. We then compute the resistance of all edges and capacitance of all nodes in the STA graph using the widely adopted -model.
@@ III-C Directed Tree Edge Weighting
and we select in our implementation. The logistic function value remains nearly 1 as approaches 1, enabling our algorithm to effectively optimize paths with poor slack. For an edge in the directed tree, its weight is defined as the maximum weight of all sinks in its downstream subtree:
@@ III-D Layer Assignment with Look Ahead Strategy
To concurrently optimize slack, power consumption, and congestion, we integrate three cost functions: net delay cost, net capacitance cost, and net congestion cost. Net capacitance is prioritized here, as it directly influences both power consumption and cell delay. In the Nangate 45nm library, cell delay dominates path delay: a cell may exhibit equivalent drive resistances exceeding 1k, whereas the resistance of a GCell edge remains below 0.01k. In a typical congestion- and via count-driven sequential layer assignment framework, we employ tree dynamic programming (DP) to optimize the objective for each net. The tree DP method proceeds bottom-up to compute the optimal solution for each candidate subtree, then top-down to backtrack the optimal solution for the entire tree. For each node, DP retains the optimal solution for each possible layer through which the path from the node to its parent traverses. However, this method is ineffective for net delay optimization due to the after-effect. Figure 8 shows the RC tree extracted from net 2 in Figure 6b using the -model, where node 3 is the root of the tree and node 5, 6, 7 are the leaves. In the Elmore delay model, the delay from node 3 to node 4 is computed as . The subtree choices for nodes 5, 6, and 7 introduce after-effects on , which is absent in congestion or wirelength cost calculations. Due to the significant resistance of vias, routing a simple 2-pin net on lower layers reduces net delay but increases net capacitance. As a result, small subtrees tend to favor lower layers for reduced delay, which can degrade the delay of upstream connections. To address this, we propose a look-ahead strategy. Let denote the cost of the subtree rooted at node comprising delay, capacitance, and congestion costs. In our layer assignment directed tree, we estimate the upstream resistance of each node and use to select optimal candidate instead of , where is the weight of delay cost, is the estimated upstream resistance of node and is the downstream capacitance of node . This approach anticipates after-effects by evaluating how downstream capacitance contributes to the entire tree’s delay cost. For example in Figure 6b and 8, we first estimate resistances and using node distances and the average unit-length metal resistance . We then use to evaluate all the candidates solution for subtree 6. Figure 9 illustrates our layer assignment flow. After the preparation steps, we obtain the directed tree structure, level information, weighting of each edge, and resistance values for the look-ahead strategy. The algorithm then proceeds to the most computation-intensive part in Algorithm 2.
Algorithm 3 details our GPU kernel implementation. Each GPU thread computes the cost and look-ahead cost of node at layer . The implementation first calculates the via delay cost for each pin at the node (Lines 4–7), then enumerates all possible via connection pairs (connect layer to layer ) used at node and its son nodes (Lines 9–26). For each son node, the candidate subtree with the minimum look-ahead cost is selected, and finally the cost, corresponding solution, and downstream load capacitance are recorded. Fig. 10 shows the enumeration process at the parent node, where a via connection from layer to is used and son nodes must connect by wires on these layers. The optimal legal candidate with minimal look-ahead cost is chosen for each son.
@@
| Case | Quality Score | Runtime (s) | ||||||||||
| 1st | Imp. | 2nd | Imp. | 3rd | Imp. | Ours | 1st | 2nd | 3rd | Ours† | GAP‡ | |
| 1 | 0.896 | -2.055 | 0.639 | -1.797 | 0.094 | -1.253 | -1.158 | 5 | 6 | 14 | 9 | 4 |
| 2 | 0.562 | -0.640 | 0.722 | -0.800 | 0.094 | -0.172 | -0.078 | 22 | 16 | 32 | 29 | 9 |
| 3 | -1.995 | -0.137 | -2.077 | -0.055 | -1.824 | -0.308 | -2.132 | 7 | 6 | 17 | 12 | 6 |
| 4 | -0.405 | +0.072 | 0.401 | -0.734 | 1.370 | -1.703 | -0.333 | 7 | 8 | 15 | 8 | 4 |
| 5 | -2.029 | -1.671 | -1.295 | -2.406 | -3.599 | -0.101 | -3.700 | 98 | 111 | 197 | 117 | 43 |
| 6 | 0.510 | -0.135 | 0.399 | -0.024 | 0.714 | -0.339 | 0.375 | 330 | 454 | 792 | 513 | 88 |
| 7 | 1.792 | -0.267 | 1.659 | -0.133 | 1.531 | -0.005 | 1.525 | 6 | 6 | 18 | 9 | 4 |
| 8 | 0.849 | -0.409 | 0.415 | +0.026 | 0.689 | -0.248 | 0.441 | 25 | 16 | 38 | 32 | 13 |
| 9 | 1.431 | -0.077 | 1.661 | -0.307 | 1.392 | -0.038 | 1.354 | 6 | 9 | 15 | 8 | 5 |
| 10 | 0.344 | -0.122 | 2.074 | -1.853 | 1.803 | -1.582 | 0.221 | 6 | 9 | 15 | 8 | 3 |
| 11 | 1.332 | -0.000 | 1.232 | +0.099 | 3.116 | -1.784 | 1.331 | 98 | 110 | 200 | 123 | 43 |
| 12 | 2.327 | -0.500 | 2.370 | -0.544 | 2.677 | -0.851 | 1.826 | 351 | 418 | 805 | 414 | 87 |
| Avg. | 0.468 | 0.495 | 0.683 | 0.711 | 0.671 | 0.699 | -0.027 | 80 | 97 | 180 | 107 | 26 |
and denote the runtime of our complete global routing flow and our proposed GAP-LA, respectively.
| Case | WNS | TNS | Power | Congestion | ||||||||||||
@@
| 1 | -0.432 | -0.398 | -0.381 | -0.365 | -1325.8 | -1340.2 | -1138.4 | -997.9 | 0.646092 | 0.646078 | 0.646080 | 0.646076 | 5846461 | 5902619 | 8003814 | 6662339 |
| 2 | -0.420 | -0.424 | -0.407 | 0.388 | -10366.2 | -10774.9 | -9344.3 | -9602.9 | 3.05441 | 3.05301 | 3.05345 | 3.05349 | 21332547 | 20468992 | 25412296 | 22947622 |
| 3 | -56.919 | -55.841 | -55.914 | -56.253 | -521344.6 | -514660.3 | -514674.3 | -515014.3 | 2.94054 | 2.94011 | 2.93806 | 2.93766 | 13315606 | 13686714 | 15687876 | 13562374 |
| 4 | -0.379 | -0.489 | -0.520 | -0.387 | -1610.0 | -2225.5 | -2524.3 | -1711.6 | 0.14310 | 0.14515 | 0.14522 | 0.14383 | 3019425 | 2478318 | 3469019 | 2689765 |
| 5 | -0.305 | -0.347 | -0.275 | -0.327 | -20473.8 | -21709.1 | -19338.7 | -17619.7 | 7.69208 | 7.73386 | 7.55296 | 7.59214 | 55005746 | 49071540 | 97506553 | 67934233 |
| 6 | -0.266 | -0.315 | -0.276 | -0.247 | -59712.5 | -62460.0 | -59533.5 | -52800.0 | 23.4497 | 23.5781 | 23.2200 | 23.1486 | 236920427 | 192012708 | 289712328 | 24407995 |
| 7 | -1.756 | -0.737 | -0.633 | -0.645 | -650.1 | -149.5 | -162.4 | -111.5 | 0.15614 | 0.15613 | 0.15611 | 0.15614 | 4349151 | 4651055 | 4386728 | 4371048 |
| 8 | -3.984 | 0.000 | -2.747 | 0.000 | -2859.8 | 0.000 | -2936.0 | 0.000 | 0.30473 | 0.30477 | 0.30478 | 0.30481 | 22567636 | 21307154 | 20564539 | 22506663 |
| 9 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.13625 | 0.13626 | 0.13628 | 0.13629 | 14061155 | 16352692 | 13637302 | 13242873 |
| 10 | -0.342 | -0.472 | -0.440 | 0.353 | -1356.1 | -1920.7 | -1806.8 | -1312.2 | 0.14303 | 0.14485 | 0.14520 | 0.14390 | 2232008 | 2122543 | 2168599 | 2087873 |
| 11 | -0.433 | -0.460 | -0.373 | -0.404 | -32016.1 | -34846.1 | -26801.4 | -27116.0 | 8.36880 | 8.35859 | 7.94192 | 8.19256 | 52783071 | 49722314 | 132027917 | 68117588 |
| 12 | -0.223 | -0.294 | -0.200 | -0.199 | -42564.1 | -50372.4 | -39516.2 | -38106.6 | 24.2439 | 24.3801 | 23.9679 | 24.0595 | 242737068 | 214110807 | 335020596 | 232170935 |

### v2

@@ Abstract
Layer assignment is critical for global routing of VLSI circuits. It converts 2D routing paths into 3D routing solutions by determining the proper metal layer for each routing segments to minimize congestion and via count. As different layers have different unit resistance and capacitance, layer assignment also has significant impacts to timing and power. With growing design complexity, it becomes increasingly challenging to simultaneously optimize timing, power, and congestion efficiently. Existing studies are mostly limited to a subset of objectives.
In this paper, we propose a GPU-accelerated performance-driven layer assignment framework, GAP-LA, for holistic optimization the aforementioned objectives. Experimental results demonstrate that we can achieve 0.3%-9.9% better worst negative slack (WNS) and 2.0%-5.4% better total negative slack (TNS) while maintaining power and congestion with competitive runtime compared with ISPD 2025 contest winners, especially on designs with up to 12 millions of nets. We also apply our method to the 2D net topology of ISPD 2025 contest winners. Compared with their original flow, we achieve 8.5%-18.3% better WNS and 2.0%-5.1% better TNS.
I Introduction
@@
Delay-driven layer assignment focuses on minimizing maximal or average delay of routing segments under congestion constraints. [31, 22] prioritize nets with large delay and formulate the delay optimization problem as an ILP problem. [8, 32, 11, 23] take coupling effect, slew and non-default-rule routing into account to better handle delay optimization problem in advanced technology node. They also apply edge weighting strategy to optimize maximal and average delay of sinks effectively. It usually takes around ten minutes for those CPU layer assignment frameworks to handle the design with around ten million pins and one million GCells.
@@ II-E Evaluation Metrics
where are input for each design to adjust the weights of different metrics and are input reference metric value. is the number of end points in the design and is total overflow of all the GCell edges. The end points usually indicate the primary outputs and D-pins of flip-flops in the design. In our evaluation flow, WNS, TNS and power are reported by open source tool OpenROAD [1]. Note that, value of does not influence the comparison among different solutions because value of is constant.
@@
as its overflow, where is for , for , and denotes the overflow weight of layer , which is given in the benchmark. The overflow metrics in Eq. 3 is better than that in Eq. 2, because it penalizes a highly congested edge in the routing solution, which makes it very difficult to resolve design rule violations in the detailed routing stage.
@@ III-A Critical Net Analysis-based Net Ordering and Batching Strategy
The delay of critical paths is of paramount importance for design timing closure, and our solution aims to minimize critical path delay as much as possible. However, since existing work often adopts net-based [14, 13, 30, 28, 32] or edge-based [2] optimization strategies for layer assignment, the delay of critical paths cannot be optimized directly. Therefore, identifying the criticality of each net is essential for our net-based approach.
Figure 5 depicts an example of a static timing analysis graph with three primary inputs and two primary outputs. Our static timing analysis engine reports one critical path for each endpoint. Critical path 1 traverses nets , and while critical path 2 traverses nets , and . We define net criticality as the number of critical paths traversing the net. In example shown in Figure 5, net exhibits the highest criticality, whereas nets and have lower criticality, and all other nets have the lowest criticality. Intuitively, optimizing a net with higher criticality yields greater benefits, as it may dominate the arrival times of multiple primary outputs.
@@
However, the strategy described above is purely sequential and must be adapted for parallel computing. Our observations indicate that nets with similar criticality and slack have comparable impacts on final design performance. As a result, strict net ordering should be relaxed: we partition the net set into multiple subsets, enabling parallel assignment within each subset with minimal quality degradation. For critical and semi-critical nets, let denote the maximum criticality. We place nets with criticality into the first subset, into the second subsets, into the third subsets and so on. For other nets with bad slack, we place nets with slack into the first subset, into the second subset, into the third subset and so on. For non-critical nets (better than WNS), we employ a traditional congestion-driven strategy instead, as these nets are far less likely to degrade design performance. Algorithm 1 outlines our critical net analysis-based net ordering and batching strategy flow. We get pin slack and critical paths from timing report offered by a static timing analysis engine and count the number of critical paths passing through each net in Line 1 and Line 2. Then the netlist is divided into three subsets, according to the timing information in Line 3. For timing-critical nets in , we partition them with the strategy described above to prioritize more critical nets in Line 4 and Line 5. In Line 6-8, we implement classical batching method used in [24, 33] to generate batches.
@@ III-B Layer Assignment Directed Tree
To facilitate the extraction of parasitic parameters and the implementation of a fine-grained parallel layer assignment algorithm, we designed a directed tree structure. In our layer assignment algorithm, we assume that a piece of 2D routing segment with no bend points should be assigned to a single layer without layer shifting, which is only performed at bends and Steiner points. To satisfy these constraints, we introduce additional nodes to ensure all inter-node connections have no bend points, so layer shifting operations are confined to nodes in the directed tree. As illustrated in Figure 6a and Figure 6b, node 6 is added to bridge node 4 and node 7, ensuring no routing paths with bend points exist in the directed tree.
Based on the tree discussed above, we extract parasitic parameters for static timing analysis. The directed tree acts as a subgraph in the STA graph, with nodes connected to relevant pin nodes in the STA graph via zero-resistance edges. For horizontal wire connections, we calculate resistance and capacitance using the average per-unit-length resistance and capacitance of all horizontal wires. For vertical wire connections, we follow the same approach. We then compute the resistance of all edges and capacitance of all nodes in the STA graph using the widely adopted -model.
@@ III-C Directed Tree Edge Weighting
and we select in our implementation.
The logistic function value remains nearly 1 as approaches 1, enabling our algorithm to effectively optimize paths with poor slack. For an edge in the directed tree, its weight is defined as the maximum weight of all sinks in its downstream subtree:
@@ III-D Layer Assignment with Look Ahead Strategy
To concurrently optimize slack, power consumption, and congestion, we integrate three cost functions: net delay cost, net capacitance cost, and net congestion cost. Net capacitance is prioritized here, as it directly influences both power consumption and cell delay. In the Nangate 45nm library, cell delay dominates path delay: a cell may exhibit equivalent drive resistances exceeding 1k, whereas the resistance of a GCell edge remains below 0.01k. In a typical congestion- and via count-driven sequential layer assignment framework, we employ tree dynamic programming (DP) to optimize the objective for each net. The tree DP method proceeds bottom-up to compute the optimal solution for each candidate subtree, then top-down to backtrack the optimal solution for the entire tree. For each node, DP retains the optimal solution for each possible layer through which the path from the node to its parent traverses. However, this method is ineffective for net delay optimization due to the after-effect.
| Layer | R (k/m) | C (fF/m) | R of Upper Via (k) |
| metal1 | |||
| metal2 | |||
| metal3 | |||
| metal4 | |||
| metal5 | |||
| metal6 | |||
| metal7 | |||
| metal8 | |||
| metal9 | |||
| metal10 | - |
Figure 8 shows the RC tree extracted from net 2 in Figure 6b using the -model, where node 3 is the root of the tree and node 5, 6, 7 are the leaves. In the Elmore delay model, the delay from node 3 to node 4 is computed as . The subtree choices for nodes 5, 6, and 7 introduce after-effects on , which is absent in congestion or wirelength cost calculations.
@@
Due to the significant resistance of vias, routing a simple 2-pin net on lower layers reduces net delay but increases net capacitance. As a result, small subtrees tend to favor lower layers for reduced delay, which can degrade the delay of upstream connections. To address this, we propose a look-ahead strategy. Let denote the cost of the subtree rooted at node comprising delay, capacitance, and congestion costs. In our layer assignment directed tree, we estimate the upstream resistance of each node and use to select optimal candidate instead of , where is the weight of delay cost, is the estimated upstream resistance of node and is the downstream capacitance of node .
This approach anticipates after-effects by evaluating how downstream capacitance contributes to the entire tree’s delay cost.
For example in Figure 6b and 8, we first estimate resistances and using node distances and the average unit-length metal resistance . We then use to evaluate all the candidates solution for subtree 6.
Figure 9 illustrates our layer assignment flow. After the preparation steps, we obtain the directed tree structure, level information, weighting of each edge, and resistance values for the look-ahead strategy. The algorithm then proceeds to the most computation-intensive part in Algorithm 2.
Algorithm 3 details our GPU kernel implementation. We design two kernels to calculate the node cost, node look-ahead cost and layers used by children nodes considering a node connects to its parent node via different layers. The first kernel is used to enumerate all possible via connection pairs connect layer to layer ) used at node and its son nodes when the node connects its parent node via layer (Line 3-6). The optimal legal candidate with minimal look-ahead cost is chosen for each son. The reason why we introduce layer is that layer determines the via delay to its children nodes. After the enumeration, we use the second kernel to calculate the best solution based on data calculated in the first kernel.
IV Experimental Results
| Quality Score | Runtime(s) | GAP Runtime(s) | Via Count | ||||||||||||||
| 1st | Imp. | 2nd | Imp. | 3rd | Imp. | Ours | 1st | 2nd | 3rd | Ours | GPU | CPU | 1st | 2nd | 3rd | Ours | |
| case1 | 0.896 | -2.055 | 0.639 | -1.797 | 0.094 | -1.253 | -1.158 | 6 | 6 | 15 | 8 | 2 | 2 | 869552 | 797909 | 1452629 | 1040032 |
| case2 | 0.562 | -0.640 | 0.722 | -0.800 | 0.094 | -0.172 | -0.078 | 28 | 21 | 30 | 27 | 6 | 11 | 5100730 | 4575171 | 6483299 | 5422990 |
| case3 | -1.995 | -0.137 | -2.077 | -0.055 | -1.824 | -0.308 | -2.132 | 8 | 6 | 17 | 9 | 4 | 3 | 1393389 | 1270319 | 2147099 | 1384821 |
| case4 | -0.405 | +0.072 | 0.401 | -0.734 | 1.370 | -1.703 | -0.333 | 7 | 8 | 16 | 6 | 2 | 3 | 1181943 | 1014414 | 1376740 | 1062075 |
| case5 | -2.029 | -1.671 | -1.295 | -2.406 | -3.599 | -0.101 | -3.700 | 115 | 134 | 192 | 113 | 24 | 46 | 23842158 | 21381686 | 32437038 | 26670038 |
| case6 | 0.510 | -0.135 | 0.399 | -0.024 | 0.714 | -0.339 | 0.375 | 388 | 476 | 687 | 421 | 58 | 139 | 89575117 | 75139223 | 82079269 | 91771657 |
| case7 | 1.792 | -0.267 | 1.659 | -0.133 | 1.531 | -0.005 | 1.525 | 7 | 7 | 16 | 7 | 3 | 3 | 830056 | 893782 | 765229 | 810382 |
| case8 | 0.849 | -0.409 | 0.415 | +0.026 | 0.689 | -0.248 | 0.441 | 30 | 16 | 38 | 31 | 8 | 10 | 5404700 | 4761540 | 4704104 | 5182861 |
| case9 | 1.431 | -0.077 | 1.661 | -0.307 | 1.392 | -0.038 | 1.354 | 8 | 6 | 18 | 8 | 3 | 3 | 1319759 | 1172952 | 1136300 | 1209985 |
| case10 | 0.344 | -0.122 | 2.074 | -1.853 | 1.803 | -1.582 | 0.221 | 7 | 8 | 16 | 6 | 2 | 3 | 1060382 | 967350 | 1079131 | 992786 |
| case11 | 1.332 | -0.000 | 1.232 | +0.099 | 3.116 | -1.784 | 1.331 | 114 | 96 | 173 | 106 | 25 | 43 | 23292577 | 20969891 | 33044068 | 26158463 |
| case12 | 2.327 | -0.500 | 2.370 | -0.544 | 2.677 | -0.851 | 1.826 | 391 | 385 | 655 | 431 | 60 | 156 | 89609278 | 79136523 | 116361040 | 91495810 |
| Avg. | 0.468 | 0.495 | 0.683 | 0.711 | 0.671 | 0.699 | -0.027 | 92 | 97 | 156 | 98 | 16 | 35 | 20289970 | 17673397 | 23588829 | 21100158 |
| WNS | TNS | Power | Congestion | |||||||||||||
@@ IV Experimental Results
| case1 | -0.432 | -0.398 | -0.381 | -0.365 | -1325.8 | -1340.2 | -1138.4 | -997.9 | 0.646092 | 0.646078 | 0.646080 | 0.646076 | 5846461 | 5902619 | 8003814 | 6662339 |
| case2 | -0.420 | -0.424 | -0.407 | 0.388 | -10366.2 | -10774.9 | -9344.3 | -9602.9 | 3.05441 | 3.05301 | 3.05345 | 3.05349 | 21332547 | 20468992 | 25412296 | 22947622 |
| case3 | -56.919 | -55.841 | -55.914 | -56.253 | -521344.6 | -514660.3 | -514674.3 | -515014.3 | 2.94054 | 2.94011 | 2.93806 | 2.93766 | 13315606 | 13686714 | 15687876 | 13562374 |
| case4 | -0.379 | -0.489 | -0.520 | -0.387 | -1610.0 | -2225.5 | -2524.3 | -1711.6 | 0.14310 | 0.14515 | 0.14522 | 0.14383 | 3019425 | 2478318 | 3469019 | 2689765 |
| case5 | -0.305 | -0.347 | -0.275 | -0.327 | -20473.8 | -21709.1 | -19338.7 | -17619.7 | 7.69208 | 7.73386 | 7.55296 | 7.59214 | 55005746 | 49071540 | 97506553 | 67934233 |
| case6 | -0.266 | -0.315 | -0.276 | -0.247 | -59712.5 | -62460.0 | -59533.5 | -52800.0 | 23.4497 | 23.5781 | 23.2200 | 23.1486 | 236920427 | 192012708 | 289712328 | 24407995 |
| case7 | -1.756 | -0.737 | -0.633 | -0.645 | -650.1 | -149.5 | -162.4 | -111.5 | 0.15614 | 0.15613 | 0.15611 | 0.15614 | 4349151 | 4651055 | 4386728 | 4371048 |
| case8 | -3.984 | 0.000 | -2.747 | 0.000 | -2859.8 | 0.000 | -2936.0 | 0.000 | 0.30473 | 0.30477 | 0.30478 | 0.30481 | 22567636 | 21307154 | 20564539 | 22506663 |
| case9 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.13625 | 0.13626 | 0.13628 | 0.13629 | 14061155 | 16352692 | 13637302 | 13242873 |
| case10 | -0.342 | -0.472 | -0.440 | 0.353 | -1356.1 | -1920.7 | -1806.8 | -1312.2 | 0.14303 | 0.14485 | 0.14520 | 0.14390 | 2232008 | 2122543 | 2168599 | 2087873 |
| case11 | -0.433 | -0.460 | -0.373 | -0.404 | -32016.1 | -34846.1 | -26801.4 | -27116.0 | 8.36880 | 8.35859 | 7.94192 | 8.19256 | 52783071 | 49722314 | 132027917 | 68117588 |
| case12 | -0.223 | -0.294 | -0.200 | -0.199 | -42564.1 | -50372.4 | -39516.2 | -38106.6 | 24.2439 | 24.3801 | 23.9679 | 24.0595 | 242737068 | 214110807 | 335020596 | 232170935 |